OpenStudy (colleen313):

3/2 divided by 1/4 =

OpenStudy (igigighjkl):

Sups, my name is Aeden and I can awnser that question of yours.

OpenStudy (colleen313):

Thanks... I need to show how to reach the answer.

OpenStudy (mathmale):

\[\frac{ \frac{ 3 }{ 2 } }{ \frac{ 1 }{ 4} }\]

OpenStudy (igigighjkl):

First you take the two numbers.

OpenStudy (igigighjkl):

Do wat he just did.

OpenStudy (igigighjkl):

...

OpenStudy (igigighjkl):

Wait a moment.

OpenStudy (colleen313):

I am confused...ok

OpenStudy (igigighjkl):

@sky060902 can help u with me. (:

OpenStudy (igigighjkl):

Darnit now I confused! @mathmale can u take the credit!

OpenStudy (igigighjkl):

Great hes off line.

OpenStudy (anonymous):

The answer wouold be 6/1 (or just 6 is fine) I used this website http://www.calculatorsoup.com/calculators/math/fractions.php

OpenStudy (anonymous):

take the second fraction and flip it and then multiply the two fractions

OpenStudy (igigighjkl):

YES I KNEW IT!!!!

OpenStudy (anonymous):

:D

RhondaSommer (rhondasommer):

I like how everyone is trying to explain it and then @southpaw says one sentence and @igigighjkl gets it ^_^

OpenStudy (igigighjkl):

LOL very funny.

OpenStudy (tennis5518):

it is 6 and i agreed with rhondasommer

OpenStudy (mathmale):

southpaw's "take the fract. in the denom., flip it and multiply " hits the nails on the head. So: i-gigi, what's your final result? Can you simplify the fraction?

OpenStudy (igigighjkl):

the awnser?

OpenStudy (igigighjkl):

Hello?

OpenStudy (igigighjkl):

The awnser is 2 2/3

OpenStudy (igigighjkl):

bye.

RhondaSommer (rhondasommer):